# Domain Knowledge Infused Generative Models for Drug Discovery Synthetic Data

### v1

@@ 5 Conclusions
The SMILES/Protein to PK/DTI model xImagand-DTI generates synthetic PK and DTI target property data that closely resembles real data in univariate and for downstream tasks. xImagand-DKI provides a solution for the challenge of sparse overlapping PK and DTI target property data, allowing researchers to generate data to tackle complex research questions and for high-throughput screening. Future work will expand xImagand-DKI to categorical PK and DTI properties, and scale to more datasets and larger model sizes. In future work we will look to explore additional reparameterization tricks for diffusion, such as discrete diffusion (Austin et al., 2021), to extend our methodology to be capable of learning and generating synthetic data following categorical and Log-logistic distributions common in drug discovery datasets.
Appendix A xImagand-DKI Model
A.1 Diffusion Model
Given samples from a data distribution $q(x_{0})$, we are interested in learning a model distribution $p_{\theta}(x_{0})$ that approximates $q(x_{0})$ and is easy to sample from. Jonathan et al. (2020) considers the following Markov chain with Gaussian transitions parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$:
$x_{t}=\sqrt{\alpha_{t}}x_{0}\ +\ \sqrt{1-\alpha_{t}}\epsilon$ (2)
We train with the simplified objective:
[…]
A.2 Pre-trained SMILES and Protein Encoders
@@
SP2PKDTI diffusion models need powerful semantic SMILE and Protein encoders to capture the complexity of arbitrary chemical and biological structure inputs. Given the sparsity and small size of PK datasets, encoders trained on specific SMILES-Pharmacokinetic or SMILES-Protein pairs are infeasible (Huang et al., 2021). Many transformer-based foundational models such as ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), SMILES-BERT (Wang et al., 2019), and MOLGPT (Bagal et al., 2021) have been pre-trained to deeply understand molecular and chemical structures and properties. Similar transformer-based foundation models such as ProtBERT (Elnaggar et al., 2020) have been pre-trained to deeply understand protein structures and properties. After pre-training, these foundational models can then be fine-tuned for various downstream molecular and protein tasks. Language models trained on a SMILES-only or protein-only corpus, significantly larger than the SMILES-Pharmacokinetic and SMILES-protein data, learn a richer and wider distribution of molecular, chemical, and protein structures.
@@
We test SMILES embeddings from ChemBERTa (Ahmad et al., 2022) and protein embeddings from ProtBERT (Elnaggar et al., 2020) trained on SMILES-only and protein-only corpora, respectively. Both embedding models were collected through the Huggingface (Wolf et al., 2020) Model Hub. Similar to Saharia et al. (2022), we freeze the weights of our embedding models. Because embeddings are computed offline, freezing the weights minimizes computation and memory footprint for embeddings during model training.
A.3 Drug Discovery Domain Knowledge
Molecular fingerprints are bit strings that encode the structural information of a molecule, such as the presence or absence of specific chemical groups, atom types, or topological features (Hu et al., 2023b). Molecular fingerprints offer a versatile representation where different algorithms tailored to capture different aspects of molecular structure, such as key-based fingerprints and hash fingerprints. Key-based fingerprints, including MACCS (Durant et al., 2002) and RDKit (Landrum, 2013), utilize a predefined fragment library to encode each molecule into a binary bit stream according to its substructure. Hash-based fingerprints such as Morgan fingerprints (Morgan, 1965) encode substructures in a molecule based on paths around atoms in a molecule. Leveraging fingerprints alongside SMILES representations in parallel increases the generalizability of models (Schimunek et al., 2023).
Appendix B Pharmacokinetic and Drug-Target Interaction Datasets
Dataset Caco. Lipo. AqSo. Free. PPBR VDss Half Cl H Cl M
DTI Overlap 338 2789 1189 112 1241 184 486 698 794
PK Overlap 179 1751 884 527 1296 163 337 879 1018
Dataset Size 906 4200 9982 642 1797 1111 665 1020 1102
All 9 PK and 3 DTI datasets are collected from TDCommons (Huang et al., 2021). Analyzing the overlap of 9 PK and 3 DTI datasets used in this study, Table 3 reveals minimal overlap and significant sparsity across datasets. We select PK datasets suitable for regression from the absorption, distribution, metabolism, and excretion (ADME) and Toxicity categories. We select DTI datasets from BindingDB (Liu et al., 2007) covering properties such as inhibition constant ($K_{i}$), dissociation constant ($K_{d}$), and half maximal inhibitory concentration (IC50). Revealing the overlap sparsity between DTI and PK, out of around 700k molecules from BindingDB, only around 5k molecules (0.7%) have PK properties defined from one of the 11 PK datasets.
The inhibition constant ($K_{i}$) is a measure of how strongly an inhibitor binds to an enzyme, effectively indicating the inhibitor’s potency. BindingDB has 375k pairs of $K_{i}$ values from 175k drugs and 3k proteins. The dissociation constant quantifies binding affinity between a drug and its target protein, defined as the free ligand concentration at which 50% of the protein binding sites are occupied at equilibrium. BindingDB has 52k pairs of $K_{d}$ values from 11k drugs and 1.5k proteins. The half maximal inhibitory concentration (IC50) is a measure of the potency of a substance in inhibiting a specific biological or biochemical function. BindingDB has 991k pairs of IC50 values from 550k drugs and 5k proteins.
Caco-2 (Wang et al., 2016) is an absorption dataset containing rates of 906 drugs passing through the Caco-2 cells, approximating the rate at which the drugs permeate through the human intestinal tissue. Lipophilicity (Wu et al., 2018) is an absorption dataset that measures the ability of 4,200 drugs to dissolve in a lipid (e.g. fats, oils) environment. AqSolDB (Sorkun et al., 2019) is an absorption dataset that measures the ability of 9,982 drugs to dissolve in water. FreeSolv (Mobley & Guthrie, 2014) is an absorption dataset that measures the experimental and calculated hydration-free energy of 642 drugs in water.
Plasma Protein Binding Rate (PPBR) (Wenlock & Tomkinson, 2016) is a distribution dataset of percentages for 1,614 drugs on how they bind to plasma proteins in the blood. Volume of Distribution at steady state (VDss) (Lombardo & Jing, 2016) is a distribution dataset that measures the degree of concentration for 1,130 drugs in body tissue compared to their concentration in blood.
Half Life (Obach et al., 2008) is an excretion dataset for 667 drugs on the duration for the concentration of the drug in the body to be reduced by half. Clearance (Di et al., 2012) is an excretion dataset for around 1,050 drugs on two clearance experiment types, microsome and hepatocyte. Drug clearance is defined as the volume of plasma cleared of a drug over a specified time (Huang et al., 2021).
Appendix C Evaluation Metrics
C.1 Hellinger Distance
$HD^{2}(p,q)=\frac{1}{2}\sum^{n}_{i=1}\left(\sqrt{p_{i}}-\sqrt{q_{i}}\right)$ (4)
C.2 Machine Learning Efficiency
Machine Learning Efficiency (MLE) is a measure that assesses the ability of the synthetic data to replicate a specific use case (Dankar & Ibrahim, 2021; Basri et al., 2023; Borisov et al., 2022). MLE represents the ability of the synthetic data to replace or augment real data in downstream use cases. To measure MLE, two models are trained separately using synthetic versus real data, and then their performance, measured by Mean-Squared Error (MSE), R-Squared (R2), and Pearson Correlation Coefficient (PCC), is evaluated on real data test sets and compared.
For this experiment, we train Linear Regression (LR) models using ChemBERTa and ProtBERT embeddings to predict each PK and DTI target property value. To prevent data leakage, we first divide real and synthetic data before combining them to form train and test sets.

### v2

@@ 6 Conclusions
The SMILES/Protein to PK/DTI model xImagand-DTI generates synthetic PK and DTI target property data that closely resembles real data in univariate and for downstream tasks. xImagand-DKI provides a solution for the challenge of sparse overlapping PK and DTI target property data, allowing researchers to generate data to tackle complex research questions and for high-throughput screening. Future work will expand xImagand-DKI to categorical PK and DTI properties, and scale to more datasets and larger model sizes. For future work, we will look to extend our model to include in vivo datasets and to investigate new applications of xImagand-DKI for QIVIVE.
Appendix A Appendix
A.1 Training Details
A.1.1 Pre-trained SMILES and Protein Encoders
@@
SP2PKDTI diffusion models need powerful semantic SMILE and protein encoders to capture the complexity of arbitrary chemical and biological structure inputs. Given the sparsity and small size of PK datasets, encoders trained on specific SMILES-Pharmacokinetic or SMILES-Protein pairs are infeasible [11]. Many transformer-based foundational models such as ChemBERTa [63, 64], SMILES-BERT [65], and MOLGPT [66] have been pre-trained to deeply understand molecular and chemical structures and properties. Similar transformer-based foundation models such as ProtBERT [67] have been pre-trained to deeply understand protein structures and properties. After pre-training, these foundational models can then be fine-tuned for various downstream molecular and protein tasks.
@@
We test SMILES embeddings from ChemBERTa [64] and protein embeddings from ProtBERT [67], trained on SMILES-only and protein-only corpora, respectively. Both embedding models were collected through the Huggingface [68] Model Hub. Similar to Saharia et al. [69], we freeze the weights of our embedding models. Because embeddings are computed offline, freezing the weights minimizes computation and memory footprint for embeddings during model training.
Imagand Model Diffusion Training
Layers 12 Learning Rate 1e-3
Heads 16 Weight Decay 5e-2
MLP Dim. 768 Epoch 3000
Emb. Dropout 10% Batch Size 256
Num Patches 48 Warmup 200
Drug Emb. Size 768 Timesteps (Train) 2000
Time Emb. Size 64 Timesteps (Infer.) 150
PK Emb. Size 256 EMA Gamma ($\gamma$) 0.994
Prot Emb. Size 1024
Drug DKI Emb. 768
Prot DKI Emb. 256
Table 4: List of Imagand Model Hyperparameters used across experiments. Model hyperparameters include the number of layers, heads, multilayered perceptron (MLP) size, embedding dropout, and sizes for the conditional, time, and pharmacokinetic (Y) embeddings. Training hyperparameters include the learning rate, weight decay, number of epochs, batch size, warmup, diffusion timesteps used for training and inference, and the Exponential Moving Average (EMA) Gamma ($\gamma$).
We train a 19M parameter model for S2PK synthesis. Model hyperparameters were not optimized and are described in Table 4. We do not find overfitting to be an issue. For classifier-free guidance, we joint-train unconditionally via dropout zeroing out sections of the SMILES embeddings with 10% probability for all of our models. For the machine learning efficiency, and univariate and bivariate distribution analysis, we utilize DeBERTa embeddings trained on PubChem and DLGN for infilling and as the noise model. We compare our model configuration to other possible configurations in the ablation experiments. All experiments were conducted using a single NVIDIA GeForce RTX 3090 GPU.
A.1.2 Static Thresholding
We apply elementwise clipping the PK predictions to [1,1][-1,1] as static thresholding, similar to Saharia et al. [69], Jonathan et al. [15]. Since PK data is min-max scaled to the same [1,1][-1,1] range as a preprocessing step, static thresholding is essential to prevent the generation of invalid and out-of-range PK values.
A.1.3 Data Processing
We first merge all 9 PK datasets to create a unified dataset containing 17k drugs over 9 unique PK columns for training and testing (90%/10% split) our models. We merge all data from 3 DTI datasets, to create a unified dataset containing 1.2M drug-protein pairs with 3 dti columns for training and testing (90%/10% split) for our models. Data from 3 DTI datasets are log-transformed. All PK and DTI data are collected from TDCommons [11].
We apply a Gaussian Quantile Transform to both PK and DTI datasets before min-max scaling between the range of [1,1][-1,1]. After removing outliers (Q11.5Q1-1.5IQR lower and Q3+1.5Q3+1.5IQR upper bound), we are left with 16.5k drugs from the original 17K drugs and 1.1M pairs from 1.2M DTI pairs. Outliers are removed to ensure that Min-Max normalization does not cause unwarranted skewness in our trainset distribution, causing issues for model training. Before infilling null values using inverse transform sampling, we store the null masks for each drug for the masked loss function.
A.1.4 Hyperparameters
A.1.5 Classifier-free Guidance
Classifier guidance uses gradients from a pre-trained model to improve quality while reducing diversity in conditional diffusion models during sampling [70]. Classifier-free guidance [71] is an alternative technique that avoids this pre-trained model by jointly training a diffusion model on conditional and unconditional objectives via dropping the condition (i.e. with 10% probability). We condition all diffusion models on learned SMILES embedding and sinusoidal time embeddings using classifier-free guidance through dropout [71, 72].
A.2 Ablation Studies
We compare Imagand with a baseline in Conditional GAN (cGAN) [48] with 1.8M parameters and Syngand [9] with 9M parameters. Similar to in Imagand, SMILES-embeddings from a pre-trained T5 model are used conditionally by the cGAN model to generate PK properties as output for a specific drug. Compared to earlier results, Table 5, Figure 7, and Figure 6 shows that Imagand is able to generate more realistic synthetic data compared to cGAN and Syngand.
Mean Std
Data Real Ours Imgd cGAN Sygd Real Ours Imgd cGAN Sygd
Caco2 0.12 0.13 0.14 0.14 0.58 0.39 0.39 0.38 0.27 0.12
Lipo 0.18 0.19 0.18 0.20 0.62 0.42 0.41 0.39 0.30 0.19
AqSol 0.11 0.18 0.11 0.13 0.10 0.41 0.39 0.36 0.29 0.18
FreeSolv 0.10 0.17 0.12 0.10 0.27 0.42 0.42 0.40 0.29 0.13
PPBR 0.57 0.59 0.56 0.63 0.97 0.50 0.49 0.47 0.37 0.09
VDss -0.60 -0.61 -0.62 -0.66 -0.98 0.44 0.43 0.39 0.31 0.06
Halflife -0.56 -0.60 -0.56 -0.61 -0.98 0.45 0.43 0.42 0.32 0.06
ClH -0.55 -0.58 -0.56 -0.59 -0.97 0.61 0.56 0.55 0.47 0.09
ClM -0.67 -0.69 -0.68 -0.74 -0.98 0.45 0.44 0.38 0.31 0.06
Table 5: Comparing mean and standard deviation values between real and synthetic target property values, rounded to two significant figures.
Figure 6: Distributions of ligand PK properties and synthetic PK Data Hellinger Distances (HDs) for cGAN. Blue, synthetic distributions; orange, real distributions.
Figure 7: Distributions of ligand PK properties (log-scale) and synthetic PK Data Hellinger Distances (HDs) for Syngand. Blue, synthetic distributions; orange, real distributions.